OpenStudy (anonymous):

You toss 3 coins. Find P(at least 1 tail is tossed).

OpenStudy (anonymous):

75%

OpenStudy (anonymous):

1 tail is 50%

OpenStudy (anonymous):

So 75% ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

for three

OpenStudy (anonymous):

for three tosses it is 75%

OpenStudy (anonymous):

Thank you sorry im just so tierd its hard to think.. ;(

OpenStudy (anonymous):

np yw

OpenStudy (anonymous):

87%??

OpenStudy (anonymous):

ok cool

OpenStudy (anonymous):

P(at least 1 tail is tossed)= 1 - p(no tail at all) p(no tail )means all three heads p(all three heads ) =1/8 hence P(at least 1 tail is tossed)= 1- 1/8 =7/8

OpenStudy (anonymous):

It was 7/8 ! Lol

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

I didn't get it matrick

OpenStudy (anonymous):

-ed

OpenStudy (anonymous):

when you toss a coin it is either heads or tails... getting a tails is 1/2 which is 50% right??? so half of 150% ( three times) is what?